# Manifolds with kinks and the asymptotic behavior of the graph Laplacian operator with Gaussian kernel

### v3

@@ 3.4.4. Riemannian distance and volume measure
The Riemannian distance defined via a length-minimizing problem extends with no change to the context of manifolds with kinks. We recall the definition for completeness and refer to [Lee, p.337–341], for instance, for more details.
@@ 3.5. Extension of Riemannian manifolds with kinks
For our purposes, we need to extend beyond the border any $\mathcal{C}^{2}$ Riemannian metric defined on a smooth manifold with kinks $M$. To this aim, we shall flow $M$ into its interior using a suitable semiflow. We adapt an argument for manifolds with corners that goes back to [DouadyHerault] at least, see also [Michor, Section 2.7].
@@ 3.5.2. Riemannian submanifolds with kinks
Let us now provide a definition of submanifold adapted to the context of manifolds with kinks, inspired by [Michor, p.19] who introduced submanifolds with corners.
@@
    Use Whitney’s theorem [Whitney] to extend each coordinate of $(\phi^{-1})^{*}g$ to form a $\mathcal{C}^{2}$ Riemannian metric $h$ on an open subset $V$ of $\mathbb{R}^{d}$ containing $\phi(x)$, such that $\phi^{-1}(V)\subset\tilde{U}$. Note that this extension may not be unique and depends on the choice of the border chart $(U,\phi)$.
@@ 4.3. Euclidean calculation
$A_{\mathcal{C}}h(0_{d}):=\int_{S^{g}\mathcal{C}}\mathop{}\!\mathrm{d}_{0_{d}}^{(2)}h\left(\theta^{(2)}\right)\mathop{}\!\mathrm{d}\sigma(\theta)\qquad\text{and}\qquad[q,h]_{\mathcal{C}}(x):=\frac{1}{2}\int_{S^{g}\mathcal{C}}\mathop{}\!\mathrm{d}_{0_{d}}h(\theta)\mathop{}\!\mathrm{d}_{0_{d}}q(\theta)\mathop{}\!\mathrm{d}\sigma(\theta).$
@@
By asymptotic property of the incomplete Gamma function (see e.g. [Pinelis])
@@
$\int_{\mathcal{C}}e^{-\|z\|^{2}}g_{1,0}(z)\mathop{}\!\mathrm{d}z$ $=q(0_{d})\int_{\mathcal{C}}e^{-\|z\|^{2}}\mathop{}\!\mathrm{d}_{0_{d}}h(z)\mathop{}\!\mathrm{d}z$
$\int_{\mathcal{C}}e^{-\|z\|^{2}}g_{1,1}(z)\mathop{}\!\mathrm{d}z$ $=\int_{\mathcal{C}}e^{-\|z\|^{2}}\mathop{}\!\mathrm{d}_{0_{d}}h(z)\mathop{}\!\mathrm{d}_{0_{d}}q(z)\mathop{}\!\mathrm{d}z$
$\int_{\mathcal{C}}e^{-\|z\|^{2}}g_{2,0}(z)\mathop{}\!\mathrm{d}z$ $=q(0_{d})\int_{\mathcal{C}}e^{-\|z\|^{2}}\mathop{}\!\mathrm{d}_{0_{d}}^{(2)}h(z^{(2)})\mathop{}\!\mathrm{d}z.$
@@ 5. Asymptotic behavior of the extrinsic Gaussian Operator
Following [CoifmanLafon2006], for a generic $y\in\mathbb{B}_{\epsilon}^{D}\cap M$ we set
@@
It follows line by line from [CoifmanLafon2006, Appendix B, Lemma 7], with the assumption that $\left\lVert y-x\right\rVert<t^{\eta}$ instead of $t^{1/2}$ like they did, which and whose implications are both indeed weaker, since $t^{\eta}>t^{1/2}$ for small $t>0$. ∎
@@
where both $O$ are independent of $\zeta\in\mathbb{R}^{d}$. Set $\alpha:=min(1,6\eta-1)$ and $\beta:=min(1,4\eta)$. Then $\alpha>0$ and $\beta>2/3$ because $\eta>1/6.$ With the above simplification and symbols, we get
@@
(36) $\begin{split}I(t)&=\int_{T^{B}_{0}(\Omega)}e^{-\left\lVert\zeta\right\rVert^{2}}\left(-\frac{1}{\sqrt{t}}\nabla\tilde{f}(0_{d})\cdot\zeta-\text{Hess}\,\tilde{f}(0_{d})(\zeta,\zeta)+o(1)\left\lVert\zeta\right\rVert^{2}\right)\\ &\quad\times(\tilde{p}(0_{d})+\sqrt{t}\nabla\tilde{p}(0_{d})\cdot\zeta+o(\sqrt{t})\left\lVert\zeta\right\rVert)(1+O(t^{\beta}))\,\mathop{}\!\mathrm{d}\zeta+o\left(\frac{1}{\sqrt{t}}\right)\\ &=\int_{T^{B}_{0}(\Omega)}e^{-\left\lVert\zeta\right\rVert^{2}}\left(-\frac{1}{\sqrt{t}}\nabla\tilde{f}(0_{d})\cdot\zeta-\text{Hess}\,\tilde{f}(0_{d})(\zeta,\zeta)+o(1)\left\lVert\zeta\right\rVert^{2}\right)\\ &\quad\times(\tilde{p}(0_{d})+\sqrt{t}\nabla\tilde{p}(0_{d})\cdot\zeta+o(\sqrt{t})\left\lVert\zeta\right\rVert)\,\mathop{}\!\mathrm{d}\zeta+O(t^{\beta-1/2})+o\left(\frac{1}{\sqrt{t}}\right)\\ &=:II(t)+O(t^{\beta-1/2}).\end{split}$
@@
(37) $\begin{split}II(t)&=-\frac{1}{\sqrt{t}}\tilde{p}(0_{d})\nabla\tilde{f}(0_{d})\cdot\int_{T^{B}_{0}(\Omega)}\zeta e^{-\left\lVert\zeta\right\rVert^{2}}\mathop{}\!\mathrm{d}\zeta-\int_{T^{B}_{0}(\Omega)}e^{-\left\lVert\zeta\right\rVert^{2}}\left(\tilde{p}(0_{d})\,\text{Hess}\,\tilde{f}(0_{d})(\zeta,\zeta)\right)\mathop{}\!\mathrm{d}\zeta\\ &\quad-\int_{T^{B}_{0}(\Omega)}e^{-\left\lVert\zeta\right\rVert^{2}}(\nabla\tilde{f}(0_{d})\cdot\zeta)(\nabla\tilde{p}(0_{d})\cdot\zeta)\mathop{}\!\mathrm{d}\zeta+O(\sqrt{t})+o\left(\frac{1}{\sqrt{t}}\right)\\ &=-\frac{c_{d}}{\sqrt{t}}(\tilde{p}(0_{d})\,\partial_{M}\tilde{f}(0_{d})+o(1))-c_{d+1}\bigg(\tilde{p}(0_{d})A_{M}\tilde{f}(0_{d})+[\tilde{p},\tilde{f}]_{M}(0_{d})\bigg)+O(\sqrt{t}),\end{split}$
@@
$\bar{L}_{t}f(x)$ $=(1+O(t^{\alpha}))(II(t)+O(t^{\beta-1/2}))$
$=II(t)[1+O(1)]+O(t^{1/6}),$
where we use that $\alpha>0$ and $\beta>1/6$ to get the second line. Together with (32) and (37), the latter implies the desired result. ∎
@@ 6.1. Interior points
If $x$ is an interior point, then $W=T_{x}M$. As a consequence,
@@
so the previous integration takes place on the empty set. Therefore, $II(t)$ and then $\mathrm{Err}(t)$ is zero for any $t>0$, as desired.
@@ 6.2. LTCDD points
$a(\theta):=\gamma^{\prime}(0_{d_{1}};\theta)\qquad\kappa(\theta):=\gamma^{\prime\prime}(0_{d-1};\theta)\,?$
@@
(39) $|\mathrm{Err}(t)|\leq 2p(x)\left(\|d_{x}f\|_{g_{x}}\frac{S_{t}}{\sqrt{t}}+\|d_{x}^{(2)}f\|_{g_{x}}\tilde{S}_{t}\right)$
where ($\triangle$ standing for the symmetric difference)
$S_{t}:=\int_{\frac{\Omega\cap\mathbb{B}_{\delta}^{d}}{\sqrt{t}}\triangle T_{0_{d-1}}^{B}\bar{\Omega}}\|\zeta\|e^{-\|\zeta\|^{2}}\mathop{}\!\mathrm{d}\zeta,\qquad\tilde{S}_{t}:=\int_{\frac{\Omega\cap\mathbb{B}_{\delta}^{d}}{\sqrt{t}}\triangle T_{0_{d-1}}^{B}\bar{\Omega}}\|\zeta\|^{2}e^{-\|\zeta\|^{2}}\mathop{}\!\mathrm{d}\zeta.$
@@
$|\mathrm{Err}(t)|$ $\leq\frac{\|p\|_{L^{\infty}(B_{t^{\eta}}(x))}\|d_{x}f\|_{g_{x}}}{\sqrt{t}}\int_{\frac{\Omega\cap\mathbb{B}_{\delta}^{d}}{\sqrt{t}}\backslash\mathbb{W}}\exp\left(-\|\zeta\|^{2}\right)\|\zeta\|d\zeta$
$+\|p\|_{L^{\infty}(B_{t^{\eta}}(x))}\Lambda(t)\int_{\frac{\Omega\cap\mathbb{B}_{\delta}^{d}}{\sqrt{t}}\backslash\mathbb{W}}\exp\left(-\|\zeta\|^{2}\right)\|\zeta\|^{2}d\xi.$
@@ 7.1. $\alpha$-Subexponential Random Variables
We begin with defining the class of $\alpha$-subexponential random variables introduced by Götze, Sambale and Sinulis in [GSS].
@@
When $\alpha=1$, the previous class coincides with classical subexponential random variables, while for $\alpha=2$, it coincides with classical subgaussian random variables, see e.g. [Vershynin] for more details about these cases. Note also that $\alpha>2$ implies subgaussian (i.e. $\alpha=2$). For values $\alpha<1$, heavier tails are permitted, as in the case of Weibull random variables, for instance.
@@
The next statement is taken from [GSS, Corollary 1.4 and Theorem 1.5]. It is a concentration result for sample means of $\alpha$-subexponential random variables.
@@
Note that (1) matches up with the subexponential tail decay given by Bernstein’s inequality in case $\alpha=1,$ while (2) coincides with the subgaussian tail decay of Hoeffding’s inequality when $\alpha=2$. We refer to [Vershynin] for a nice account on these two classical inequalities.

### v4

@@ 3.4.4. Riemannian distance and volume measure
The Riemannian distance defined via a length-minimizing problem extends with no change to the context of manifolds with kinks. We recall the definition for completeness and refer to [21, p.337–341], for instance, for more details.
@@ 3.5. Extension of Riemannian manifolds with kinks
For our purposes, we need to extend beyond the border any $\mathcal{C}^{2}$ Riemannian metric defined on a smooth manifold with kinks $M$. To this aim, we shall flow $M$ into its interior using a suitable semiflow. We adapt an argument for manifolds with corners that goes back to [11] at least, see also [24, Section 2.7].
@@ 3.5.2. Riemannian submanifolds with kinks
Let us now provide a definition of submanifold adapted to the context of manifolds with kinks, inspired by [24, p.19] who introduced submanifolds with corners.
@@
    Use Whitney’s theorem [34] to extend each coordinate of $(\phi^{-1})^{*}g$ to form a $\mathcal{C}^{2}$ Riemannian metric $h$ on an open subset $V$ of $\mathbb{R}^{d}$ containing $\phi(x)$, such that $\phi^{-1}(V)\subset\tilde{U}$. Note that this extension may not be unique and depends on the choice of the border chart $(U,\phi)$.
@@ 4.3. Euclidean calculation
$A_{\mathcal{C}}h(0_{d}):=\frac{1}{2}\int_{S^{g}\mathcal{C}}\mathop{}\!\mathrm{d}_{0_{d}}^{(2)}h\left(\theta^{(2)}\right)\mathop{}\!\mathrm{d}\sigma(\theta)\qquad\text{and}\qquad[q,h]_{\mathcal{C}}(x):=\int_{S^{g}\mathcal{C}}\mathop{}\!\mathrm{d}_{0_{d}}h(\theta)\mathop{}\!\mathrm{d}_{0_{d}}q(\theta)\mathop{}\!\mathrm{d}\sigma(\theta).$
@@
By asymptotic property of the incomplete Gamma function (see e.g. [27])
@@
$\int_{\mathcal{C}}e^{-\|z\|^{2}}h_{1,0}(z)\mathop{}\!\mathrm{d}z$ $=q(0_{d})\int_{\mathcal{C}}e^{-\|z\|^{2}}\mathop{}\!\mathrm{d}_{0_{d}}h(z)\mathop{}\!\mathrm{d}z$
$\int_{\mathcal{C}}e^{-\|z\|^{2}}h_{1,1}(z)\mathop{}\!\mathrm{d}z$ $=\int_{\mathcal{C}}e^{-\|z\|^{2}}\mathop{}\!\mathrm{d}_{0_{d}}h(z)\mathop{}\!\mathrm{d}_{0_{d}}q(z)\mathop{}\!\mathrm{d}z$
$\int_{\mathcal{C}}e^{-\|z\|^{2}}h_{2,0}(z)\mathop{}\!\mathrm{d}z$ $=q(0_{d})\int_{\mathcal{C}}e^{-\|z\|^{2}}\mathop{}\!\mathrm{d}_{0_{d}}^{(2)}h(z^{(2)})\mathop{}\!\mathrm{d}z.$
@@ 5. Asymptotic behavior of the extrinsic Gaussian Operator
Following [9], for a generic $y\in\mathbb{B}_{\epsilon}^{D}\cap M$ we set
@@
It follows line by line from [9, Appendix B, Lemma 7], with the assumption that $\left\lVert y-x\right\rVert<t^{\eta}$ instead of $t^{1/2}$ like they did, which and whose implications are both indeed weaker, since $t^{\eta}>t^{1/2}$ for small $t>0$. ∎
@@
where both $O$ are independent of $\zeta\in\mathbb{R}^{d}$. Set $\alpha:=\min(1,6\eta-1)$ and $\beta:=\min(1,4\eta)$. Then $\alpha>0$ and $\beta>2/3$ because $\eta>1/6.$ With the above simplification and symbols, we get
@@
(37) $\begin{split}I(t)&=\int_{T^{B}_{0}(\Omega)}e^{-\left\lVert\zeta\right\rVert^{2}}\left(-\frac{1}{\sqrt{t}}\nabla\tilde{f}(0_{d})\cdot\zeta-\frac{1}{2}\text{Hess}\,\tilde{f}(0_{d})(\zeta,\zeta)+o(1)\left\lVert\zeta\right\rVert^{2}\right)\\ &\quad\times(\tilde{p}(0_{d})+\sqrt{t}\nabla\tilde{p}(0_{d})\cdot\zeta+o(\sqrt{t})\left\lVert\zeta\right\rVert)(1+O(t^{\beta}))\,\mathop{}\!\mathrm{d}\zeta+o\left(\frac{1}{\sqrt{t}}\right)\\ &=\int_{T^{B}_{0}(\Omega)}e^{-\left\lVert\zeta\right\rVert^{2}}\left(-\frac{1}{\sqrt{t}}\nabla\tilde{f}(0_{d})\cdot\zeta-\frac{1}{2}\text{Hess}\,\tilde{f}(0_{d})(\zeta,\zeta)+o(1)\left\lVert\zeta\right\rVert^{2}\right)\\ &\quad\times(\tilde{p}(0_{d})+\sqrt{t}\nabla\tilde{p}(0_{d})\cdot\zeta+o(\sqrt{t})\left\lVert\zeta\right\rVert)\,\mathop{}\!\mathrm{d}\zeta+O(t^{\beta-1/2})+o\left(\frac{1}{\sqrt{t}}\right)\\ &=:II(t)+O(t^{\beta-1/2}).\end{split}$
@@
(38) $\begin{split}II(t)&=-\frac{1}{\sqrt{t}}\tilde{p}(0_{d})\nabla\tilde{f}(0_{d})\cdot\int_{T^{B}_{0}(\Omega)}\zeta e^{-\left\lVert\zeta\right\rVert^{2}}\mathop{}\!\mathrm{d}\zeta-\frac{1}{2}\int_{T^{B}_{0}(\Omega)}e^{-\left\lVert\zeta\right\rVert^{2}}\tilde{p}(0_{d})\,\text{Hess}\,\tilde{f}(0_{d})(\zeta,\zeta)\mathop{}\!\mathrm{d}\zeta\\ &\quad-\int_{T^{B}_{0}(\Omega)}e^{-\left\lVert\zeta\right\rVert^{2}}(\nabla\tilde{f}(0_{d})\cdot\zeta)(\nabla\tilde{p}(0_{d})\cdot\zeta)\mathop{}\!\mathrm{d}\zeta+O(\sqrt{t})+o\left(\frac{1}{\sqrt{t}}\right)\\ &=-\frac{c_{d}}{\sqrt{t}}(\tilde{p}(0_{d})\,\partial_{M}\tilde{f}(0_{d})+o(1))-c_{d+1}\bigg(\tilde{p}(0_{d})A_{M}\tilde{f}(0_{d})+[\tilde{p},\tilde{f}]_{M}(0_{d})\bigg)+O(\sqrt{t}),\end{split}$
@@
$\bar{L}_{t}f(x)$ $=(1+O(t^{\alpha}))(II(t)+O(t^{\beta-1/2}))=II(t)[1+o(1)]+O(t^{1/6}),$
where we use that $\alpha>0$ and $\beta>1/6$ to get the second equality. Together with (33) and (38), the latter implies the desired result. ∎
@@ 6.1. Interior points
Assume that $x$ is an interior point. Then $I_{x}M=T_{x}M$ so $v_{g}(x)=0$ by symmetry. Moreover, $W=T_{x}M$, so
@@
then the previous integration takes place on the empty set. Therefore, $II(t)$ and then $\mathrm{Err}(t)$ is zero for any small enough $t>0$. Thus
$L_{t}f(x)$ $=\frac{1}{2}\Gamma\left(\frac{d}{2}+1\right)\bigg(p(x)A_{g}f(x)+[p,f]_{g}(x)\bigg)+O(\sqrt{t})+O(t^{-d}e^{-t^{2\eta-1}})$
as $t\downarrow 0$. That
$\frac{1}{2}\Gamma\left(\frac{d}{2}+1\right)\bigg(p(x)A_{g}f(x)+[p,f]_{g}(x)\bigg)=-\frac{\pi^{d/2}}{4}\,\Delta_{M,p}f(x)$
is a consequence of a direct calculation involving the second moments (see e.g. [25, p. 33–34])
$\int_{S^{g}_{x}M}\theta_{i}\theta_{j}\sigma(\theta)=\frac{\delta_{ij}\mathcal{H}^{d-1}(\mathbb{S}^{d-1})}{d}\,\cdot$
6.2. $\mathcal{C}^{1}$ boundary points
If $x$ is a $\mathcal{C}^{1}$ boundary point, we obtain
$I_{x}M=\{v\in T_{x}M:g(\partial_{\nu}f(x),v)\geq 0\}$
where $\partial_{\nu}f(x)$ is the inner normal derivative of $f$ at $x$, hence $\partial_{v_{g}(x)}f(x)$ coincides with $\partial_{\nu}f(x)$. Then we act as in the previous subsection to get the right constant in (3)
@@ 6.3. LTCDD points
$a(\theta):=\gamma^{\prime}(0_{d_{1}};\theta),\qquad\kappa(\theta):=\gamma^{\prime\prime}(0_{d-1};\theta).$
@@
(40) $|\mathrm{Err}(t)|\leq 2p(x)\left(\|d_{x}f\|_{g_{x}}S_{t}+\sqrt{t}\|d_{x}^{(2)}f\|_{g_{x}}\tilde{S}_{t}\right)$
where
$S_{t}:=\int_{\frac{\Omega\cap\mathbb{B}_{\delta}^{d}}{\sqrt{t}}\triangle T_{0_{d-1}}^{B}\bar{\Omega}}\|\zeta\|e^{-\|\zeta\|^{2}}\mathop{}\!\mathrm{d}\zeta,\qquad\tilde{S}_{t}:=\int_{\frac{\Omega\cap\mathbb{B}_{\delta}^{d}}{\sqrt{t}}\triangle T_{0_{d-1}}^{B}\bar{\Omega}}\|\zeta\|^{2}e^{-\|\zeta\|^{2}}\mathop{}\!\mathrm{d}\zeta,$
and $\triangle$ stands for the symmetric difference.
@@
$|\mathrm{Err}(t)|$ $\leq\|p\|_{L^{\infty}(B_{t^{\eta}}(x))}\|d_{x}f\|_{g_{x}}\int_{\frac{\Omega\cap\mathbb{B}_{\delta}^{d}}{\sqrt{t}}\backslash\mathbb{W}}\exp\left(-\|\zeta\|^{2}\right)\|\zeta\|d\zeta$
$+\sqrt{t}\|p\|_{L^{\infty}(B_{t^{\eta}}(x))}\Lambda(t)\int_{\frac{\Omega\cap\mathbb{B}_{\delta}^{d}}{\sqrt{t}}\backslash\mathbb{W}}\exp\left(-\|\zeta\|^{2}\right)\|\zeta\|^{2}d\xi.$
@@ 7.1. $\alpha$-Subexponential Random Variables
We begin with defining the class of $\alpha$-subexponential random variables introduced by Götze, Sambale and Sinulis in [15].
@@
When $\alpha=1$, the previous class coincides with classical subexponential random variables, while for $\alpha=2$, it coincides with classical subgaussian random variables, see e.g. [33] for more details about these cases. Note also that $\alpha>2$ implies subgaussian (i.e. $\alpha=2$). For values $\alpha<1$, heavier tails are permitted, as in the case of Weibull random variables, for instance.
@@
The next statement is taken from [15, Corollary 1.4 and Theorem 1.5]. It is a concentration result for sample means of $\alpha$-subexponential random variables.
@@
Note that (1) matches up with the subexponential tail decay given by Bernstein’s inequality in case $\alpha=1,$ while (2) coincides with the subgaussian tail decay of Hoeffding’s inequality when $\alpha=2$. We refer to [33] for a nice account on these two classical inequalities.
@@ 8.3. Experiments for a 2D cusp
9. Acknowledgments.
The authors are funded by the Research Foundation – Flanders (FWO) via the Odysseus II programme no. G0DBZ23N. The first author has also been partially supported by MBZUAI SU Fund for this work. Both authors thank Dominic Joyce for helpful answers concerning manifolds with corners, Iosif Pinelis who suggested the proof of (8), and Laurent Bessières for a comment on the first version of the paper.